Herschel Status Report - March 2011
21 March 2011 11:20
Report for period 23 February to 16 March 2011
Mission operations of the Herschel space observatory continued nominally during the reporting period, with the spacecraft and subsystems all performing as expected.
The spacecraft continues to be in good health and is operating nominally.
Instrument operations for PACS, SPIRE and HIFI have been nominal for the entire reporting period, apart from a single event with HIFI.
On 28 February the memory of HIFI's Local Oscillator Control Unit (LCU) was affected in a critical area that is needed for accepting commands, providing telemetry, and for patching defective memory locations. This caused the communication to be lost between the LCU and the Instrument Control Unit. The fact that the LCU was drawing nominal power at the time suggested that it was a single event upset, or bit flip problem.
The only way in which normal instrument operations can be restored from this condition is to power cycle the LCU, which had been powered on continuously since 10 January 2010 when the redundant HIFI signal chain was put into operations. In the worst possible case the resulting power spike on the secondary converter side could have led to the loss of a diode (as had happened when on 2 August 2009 the primary chain was inadvertently power cycled - see the October 2009 status report), and thus to a loss of the instrument. Fortunately the recovery of the instrument proceeded flawlessly and HIFI has resumed nominal operations. The event requires two days of HIFI observations to be rescheduled.
Ground Segment operations have been nominal and 100% of the data continue to be recovered. As of 4 March 2011, the completion of the different programme parts was:
For more details of these different programme parts, see the "overview of Herschel observing" linked from the right-hand menu.